The sum of the digits of the two-digit number is 8. The number with the digits interchanged is 14 more than 8 times the units digit of the original number. Find the original number.
Answer by mananth(16949)

The sum of the digits of the two-digit number is 8. The number with the digits interchanged is 14 more than 8 times the units digit of the original number. Find the original number.
..
let ten's digit be x
and units digit be y
x+y=8..................1
the number = 10x+y
..
10y+x = 14+8y
10y-8y+x=14
2y+x=14................2
..
subtract equation 2 from equation1
(x+y)-(2y+x)=8-14
x+y-2y-x=-6
-y=-6
y=6
..
Plug the value of y in equation 1
x+6=8
x=2
The number = 26